OpenStudy (littleheron):

Solve for y. 2y -10 = -4x y = 2x - 5 y = -2x + 5 y = 2x + 5 y = -2x - 5

OpenStudy (jjamz87):

B

OpenStudy (mathmale):

To "Solve for y. 2y -10 = -4x" you need to isolate y on the left side of your equation. How would you get rid of the -10 on the left side? Hint: You must perform a single operation on both sides of the equation.

OpenStudy (18turns):

hes making him do the work to figure it out

OpenStudy (18turns):

so he learns how to do it

OpenStudy (littleheron):

im a girl.,...

OpenStudy (jjamz87):

(S)He's right. sorry. I'm not one for writing thinout

OpenStudy (18turns):

sorry mam

OpenStudy (just_one_last_goodbye):

@jjamz87 Please explain your answer ^_^ direct answers are not allowed due to the asker is not learning, so please lead the user to the answer c: when you signed up you agreed to these rules. Thank you. http://openstudy.com/code-of-conduct

OpenStudy (jjamz87):

ok. Solve for y: 2y - 10 = -4x divide both sides by 2 y - 5 = -2x add 5 to both sides y = -2x + 5 There's your answer

OpenStudy (jjamz87):

I apologize.

OpenStudy (just_one_last_goodbye):

Thank you ^_^